### v1

# Large Language Models for Multilingual Vulnerability Detection: How Far Are We?

Abstract.
Various deep learning-based approaches utilizing pre-trained language models (PLMs) have been proposed for automated vulnerability detection. With recent advancements in large language models (LLMs), several studies have begun exploring their application to vulnerability detection tasks. However, existing studies primarily focus on specific programming languages (e.g., C/C++) and function-level detection, leaving the strengths and weaknesses of PLMs and LLMs in multilingual and multi-granularity scenarios largely unexplored. To bridge this gap, we conduct a comprehensive fine-grained empirical study evaluating the effectiveness of state-of-the-art PLMs and LLMs for multilingual vulnerability detection. Using over 30,000 real-world vulnerability-fixing patches across seven programming languages, we systematically assess model performance at both the function-level and line-level. Our key findings indicate that GPT-4o, enhanced through instruction tuning and few-shot prompting, significantly outperforms all other evaluated models, including CodeT5P. Furthermore, the LLM-based approach demonstrates superior capability in detecting unique multilingual vulnerabilities, particularly excelling in identifying the most dangerous and high-severity vulnerabilities. These results underscore the promising potential of adopting LLMs for multilingual vulnerability detection at function-level and line-level, revealing their complementary strengths and substantial improvements over PLM approaches. This first empirical evaluation of PLMs and LLMs for multilingual vulnerability detection highlights LLMs’ value in addressing real-world software security challenges.
Multilingual Vulnerability, Vulnerability Detection, Large Language Model
journal: JACMarticle: 111copyright: acmcopyrightccs: Software and its engineering
1. Introduction
Software vulnerabilities are flaws in code that pose significant risks to software systems, potentially leading to compromised sensitive information and system failures. To address these risks, researchers have developed various automated vulnerability detection (AVD) techniques. These techniques typically fall into three main categories: rule-based techniques, machine learning (ML)-based techniques, and deep learning (DL)-based techniques. Rule-based techniques (Gobbi and Kinder, 2023; Kim et al., 2019) and ML-based techniques (Grieco et al., 2016; Scandariato et al., 2014) require significant human expertise to design rules and features, yet often struggle to define effective patterns for identifying vulnerabilities. DL-based techniques often rely on Recurrent Neural Networks (RNNs) (Li et al., 2021b, a; Zou et al., 2021), Graph Neural Networks (GNNs) (Chakraborty et al., 2021; Zhou et al., 2019; Nguyen et al., 2022; Steenhoek et al., 2024), and Pre-trained Language Models (PLMs) (Hanif and Maffeis, 2022; Fu and Tantithamthavorn, 2022; Liu et al., 2024). These methods are relatively effective since they can learn vulnerability patterns in an end-to-end manner. Particularly, leveraging pre-trained knowledge, PLMs formulate vulnerability detection as a binary classification on code embedding, proven superior to RNNs and GNNs at detecting vulnerabilities (Steenhoek et al., 2023). Although PLMs show promise to some extent, these methods still face limitations in fully understanding code semantics for extracting representative or unseen vulnerability patterns (Ding et al., 2024).
Large Language Models (LLMs) have recently garnered widespread attention and shown strong potential across a variety of code-related tasks (Hou et al., 2023). Several empirical studies (Fu et al., 2023a; Zhou et al., 2024a; Yin et al., 2024; Yin, 2024; Zhou et al., 2024c) have explored using LLMs for vulnerability detection. For instance, Fu et al. (2023a) found that GPT-3.5-Turbo and GPT-4 fail to detect single-language (i.e., C/C++) vulnerabilities at both function-level and line-level.  Yin (2024) revealed that ChatGPT, when using prompts alone, can be easily swayed to change vulnerability classifications, indicating low confidence. Furthermore,  Yin et al. (2024) demonstrated that while fine-tuned LLMs can detect vulnerabilities, they perform more weakly than PLMs. Moreover, some LLM-based AVD techniques have been proposed, including Vul-RAG (Du et al., 2024), MSIVD (Yang et al., 2024), and GRACE (Lu et al., 2024). These techniques demonstrate that LLMs are competitive with PLMs at detecting vulnerabilities in individual programming languages. For example, MSIVD applied multitask learning with LoRA to fine-tune the LLM and incorporated fine-tuned LLM embedding and Abstract Syntax Tree (AST) embedding. Vul-RAG checks code vulnerability by reasoning about vulnerability causes and fixing solutions from retrieved vulnerability knowledge, proving LLMs’ potential in detecting code vulnerabilities.
While LLMs-based and PLMs-based techniques have made significant progress on vulnerability detection, we identify three remaining limitations: (I) Focus on a single programming language. Most existing studies limit their vulnerability detection capabilities to C and C++, often relying on datasets such as CVEfixes (Bhandari et al., 2021) and Big-Vul (Fan et al., 2020). However, real-world development environments are inherently complex and often involve multiple programming languages simultaneously (Li et al., 2022), with existing studies further revealing the prevalence of vulnerabilities across diverse language ecosystems such as Python (Alfadel et al., 2023) and Go (Hu et al., 2024). To be practically applicable, research on AVD needs to expand beyond single-language vulnerability to address multilingual vulnerability. (II) Unclear comparative strengths of PLMs and LLMs on multilingual vulnerability scenario. PLMs and LLMs are considered two promising mainstream approaches. The relative advantages and limitations of PLMs versus LLMs in multilingual vulnerability detection remain underexplored. Additionally, existing LLM-based approaches often overlook the diversity of available LLMs and the variety of applicable prompting and fine-tuning strategies. This knowledge gap in understanding makes it difficult to select and improve models for specific tasks in multilingual contexts. (III) Lack of comprehensive evaluation across granularity levels. Vulnerability detection typically operates at two levels: function-level and line-level. Function-level detection determines whether a given function contains any vulnerabilities. However, the coarse granularity of this approach can hinder developers from efficiently identifying and addressing specific issues. In contrast, line-level detection aims to pinpoint the exact lines within a vulnerable function that are responsible for the vulnerability. Existing research (Zhou et al., 2024a) has primarily focused on function-level vulnerability detection in single-language settings, whereas the investigation of line-level multilingual vulnerability detection remains significantly more pressing. Additionally, multilingual contexts pose greater challenges for line-level vulnerability detection, as variations in syntax and semantics across programming languages increase the difficulty for language models to generalize effectively.
In this work, we conduct an empirical study to systematically investigate the effectiveness of existing PLMs and examine the performance of advanced LLMs in multilingual vulnerability detection. Our hypothesis is that LLMs, with their remarkable semantic understanding and language-agnostic capabilities, combined with effective learning strategies, can enhance the effectiveness of detecting function-level and line-level multilingual vulnerability. For our evaluation, we utilize REEF, a comprehensive multilingual vulnerability corpus containing 4,466 CVEs with 30,987 patches. This dataset spans seven major programming languages: C, C#, C++, Go, JavaScript, Java, and Python. To structure our investigation, we propose the following three research questions:
- RQ1: How effective are PLMs and LLMs in detecting multilingual vulnerabilities at the function level?
- RQ2: How effective are PLMs and LLMs in detecting multilingual vulnerabilities at the line level?
- RQ3: What are the strengths and weaknesses of the PLMs and LLMs in multilingual vulnerability detection?
Our key empirical findings are: (1) An appropriate LLM with instruction tuning and few-shot prompting demonstrates promising multilingual vulnerability detection performance. Specifically, for both function-level and line-level multilingual vulnerability detection, GPT-4o with instruction tuning and few-shot prompting demonstrates superior performance, achieving the highest accuracy (0.7196) at function-level compared to the best-performing CodeT5P (0.6037), and the best F1-score (0.6641) at line-level compared to other studied PLMs and LLMs. (2) Across different programming languages, the best-performing LLM achieves its highest function-level accuracy (0.8082) with Go and its lowest (0.6626) with Python. At the line-level, GPT-4o attains its highest F1-score (0.7815) on JavaScript, while its lowest F1-score (0.4348) is observed on C#. (3) Orthogonality analysis reveals that the best-performing LLM excels not only in uniquely correct detections but also shows higher accuracy in identifying the top-25 most dangerous CWE-IDs compared to PLMs. Base on CVSS severity, the top LLM also surpasses other studied PLMs and LLMs in detecting high-severity vulnerabilities (e.g., Critical and High level severity). (4) Furthermore, we investigate the impact of LLM size and compare effectiveness between reasoning and non-reasoning LLMs. The results show that the size of LLMs is not the decisive factor affecting performance on multilingual vulnerability detection, and the use of reasoning LLMs does not lead to significant improvements.
Contributions. To sum up, the contributions of this study are:
  1. (1)
    This is the first empirical study to comprehensively investigate the effectiveness of PLMs and LLMs with various strategies on both function-level and line-level multilingual vulnerability detection.
  2. (2)
    The empirical results confirm the promising role of LLMs in multilingual vulnerability detection at function-level and line-level perspectives, particularly when instruction-tuning LLMs with few-shot prompting strategies.
  3. (3)
    Through fine-grained analysis, we provide valuable insights into the capabilities and limitations of LLMs for multilingual vulnerability detection, offering essential guidance for future research aimed at advancing LLM-based vulnerability detection techniques.
  4. (4)
    We have made the replication package publicly available on our homepage (rep, [n. d.]), including all used data, codes, and analysis details involved in our study.
Paper Extension. This paper extends our prior work on a preliminary study (Yu et al., 2025), published as a research paper at the 1st International Workshop on Large Language Model Supply Chain Analysis, co-located with ISSTA’25. The key differences can be summarized as follows: (I) A broader range of studied models and prompting strategies. Specifically, we enhance PLMs by incorporating three OpenAI text-embedding models, and for LLMs, we explore two additional strategies: few-shot prompting and instruction tuning. (II) Incorporation of line-level detection. To provide a more comprehensive perspective, we extend the research scope beyond function-level detection to include line-level detection, enabling more detailed analysis of vulnerability locations. (III) Finer-grained analyses of detection performance. We conduct an in-depth analysis focusing on unique correct and incorrect detections, prediction tendencies for the Top-25 most dangerous CWE-IDs, and the effectiveness of detecting vulnerabilities across different severity levels. (IV) Exploration of effects of model size and architecture. We further investigate how the size of open-source LLMs and the use of reasoning-capable LLMs impact performance in multilingual vulnerability detection.
2. Related Work
2.1. Software Vulnerability
Software vulnerabilities are security flaws in software systems that malicious actors can exploit. These vulnerabilities can cause significant damage when left unpatched, particularly in widely-used software, leading to substantial financial losses (Bilge and Dumitras, [n. d.]). For example, the Log4Shell vulnerability (CVE-2021-44228) (Luttwak and Schindel, 2021) in the Log4j library, which allowed attackers to execute arbitrary Java code on servers and other systems, exposed approximately 93% of enterprise cloud environments to significant security risks. Two main classification systems in the community help categorize these issues: Common Weakness Enumeration (CWE) for identifying weakness types, and Common Vulnerability Exposure (CVE) for documenting specific vulnerability instances (CWE, 2024; CVE, 2024).
Recent studies highlight the prevalence of vulnerabilities in open-source software.  Jia et al. (2022) found that the Cargo ecosystem’s security vulnerabilities are primarily memory-related, with 18% of affected libraries still vulnerable in their latest versions, and 19.78% of all versions in the ecosystem impacted by vulnerability propagation.  Zerouali et al. (2022) observed that vulnerabilities in npm packages affect a median of 30 package releases, compared to 59 releases in RubyGems packages. Furthermore, a significant proportion of external GitHub projects is exposed to vulnerabilities originating from both direct and indirect dependencies.  Li et al. (2022) presented a large-scale study of popular multilingual projects on GitHub and identified statistically significant associations between the vulnerability proneness of multilingual code (both overall and for specific categories) and its language selection. Alfadel et al. (2023) analyzed 550 vulnerability reports related to 252 Python packages, providing insights into common security issues within the Python ecosystem. A comprehensive analysis (Mir et al., 2023) of 3 million Maven packages demonstrated that about one-third of the packages in the dataset are identified as vulnerable only when all the transitive dependencies are taken into consideration.  Hu et al. (2024) investigated the prevalence and remediation delays of vulnerabilities in Go modules, finding that 66.10% of modules are affected and identifying two types of delays that impede the timely resolution of vulnerabilities. These empirical findings emphasize the widespread presence of vulnerabilities in open-source software across diverse programming languages, underscoring the urgent need for effective vulnerability detection in multilingual contexts.
2.2. Automated Vulnerability Detection
Automated vulnerability detection is a security assessment method that employs techniques such as static code analysis (Gobbi and Kinder, 2023) or dynamic execution testing (Kim et al., 2019) to automatically identify potential security weaknesses in software or systems. Recent breakthroughs in deep learning, especially PLMs, have revolutionized automated vulnerability detection, leading to significant advances in cybersecurity applications (Zhou et al., 2024a; Steenhoek et al., 2023). These methods learn potential vulnerability patterns by constructing abstract representations of source code and establishing a nonlinear mapping relationship between source code characterization and vulnerability presence (i.e., whether a given code snippet contains security vulnerabilities).
Specifically,  Hanif and Maffeis (2022) introduced VulBERTa, a deep learning model that leverages a custom tokenization pipeline to pre-train RoBERTa (Liu et al., 2019) on real-world C/C++ code, enabling enhanced vulnerability detection through learned code syntax and semantics. LineVul (Fu and Tantithamthavorn, 2022) employs a BERT-based architecture (Devlin et al., 2019) to create code representations for detecting vulnerabilities at the function level and uses its attention mechanism to identify specific vulnerable lines. SVulD (Ni et al., 2023) trains a model to differentiate semantic representations of functions, irrespective of lexical similarity, for vulnerability detection, and provides natural language explanations to help developers intuitively understand the root causes of vulnerabilities.  Zhang et al. (2023b) decomposed the syntax-based Control Flow Graph (CFG) of a code snippet into multiple execution paths, extracted their representations through LLMs with intra- and inter-path attention mechanisms, and aggregated these representations to detect vulnerabilities.  Liu et al. (2024) pre-trains their PLM, PDBERT, by leveraging the Abstract Syntax Tree (AST) and Program Dependency Graph (PDG) of functions to predict statement-level control dependencies and token-level data dependencies. They then fine-tune the PLM for the vulnerability detection task. DeepDFA (Steenhoek et al., 2024) introduces a graph-based learning framework that incorporates graph embeddings with PLMs to enhance vulnerability detection performance. In this study, they use general PLMs as vulnerability detection baselines.
However, PLMs require vast amounts of data to uncover patterns of vulnerability, and their generalization ability may be limited when confronted with unseen data. Recently, LLMs that are pre-trained on extensive code corpora have demonstrated a remarkable ability to understand code semantics, showing significant performance improvements in code-related tasks (Hou et al., 2023). Several empirical studies (Fu et al., 2023a; Zhou et al., 2024a; Yin et al., 2024; Yin, 2024; Zhou et al., 2024c) have explored using LLMs for vulnerability detection.  Fu et al. (2023a) examined ChatGPT’s capability to detect vulnerabilities in C/C++ functions through a zero-shot prompt. They found that GPT-3.5-Turbo and GPT-4 fail to detect single-language vulnerabilities at both function-level and line-level.  Yin (2024) demonstrated that when relying solely on prompts, ChatGPT is highly susceptible to altering vulnerability classifications, reflecting low confidence in its assessments. Although the inclusion of contextual information enhances its accuracy, the model still struggles to reliably predict severity ratings for certain categories of CWEs. Vul-RAG (Du et al., 2024), a technique for vulnerability detection based on LLMs, constructs a comprehensive knowledge base of vulnerabilities, retrieves relevant information through semantic similarity, and utilizes reasoning to assess the presence of vulnerabilities in code. MSIVD (Yang et al., 2024) integrates a multitask sequence-to-sequence language model with program control flow graphs, which are encoded as graph neural networks, to enable sequence-to-classification vulnerability detection. GRACE (Lu et al., 2024) leverages graph structural information in the code and in-context learning to enhance the effectiveness of LLM-based software vulnerability detection.
Notably, researchers have observed significant variations in the detection difficulty of different vulnerability types. Studies by  Russell et al. (2018),  Zou et al. (2021), and  Xu et al. (2022) indicate that detecting certain types of vulnerabilities is more challenging. Furthermore,  Hin et al. (2022) evaluated model performance in cross-project settings using a leave-one-out approach and found a slight decline in performance under such scenarios. However, as modern software development increasingly trends toward multilingual integration, the diversity of codebases introduces new challenges for software security. The effectiveness of existing vulnerability detection methods in multilingual scenarios remains unverified, making it difficult to meet the demands of real-world development.
3. Study Design
This section introduces the overview of our study design. Initially, based on the latest multilingual vulnerability dataset, we construct the training and test datasets for the detection of multilingual vulnerability at the function and line levels. We first investigate the effectiveness of existing PLM-based AVD approaches and LLMs in detecting function-level multilingual vulnerability (RQ1). After that, we further examine the performance of existing PLM-based AVD approaches and LLMs in detecting line-level multilingual vulnerability (RQ2). Finally, we obtain an understanding of the strengths and weaknesses of the PLM-based AVD approaches and LLMs in multilingual vulnerability detection from different perspectives (RQ3).
3.1. Dataset Preparation
Studied dataset. To evaluate PLM-based AVD approaches and LLMs for multilingual vulnerability detection, we utilize the REEF dataset (Wang et al., 2023b). This dataset contains 4,466 CVEs with 30,987 patches across seven programming languages, including comprehensive vulnerability information (CVE, CWE, CVSS, etc.) and project details (such as commit messages). REEF is constructed from real-world vulnerabilities sourced from the National Vulnerability Database (NVD) and Mend’s CVE list (WhiteSource, 2022), an open-source vulnerability database covering 2016-2023. At the function-level, the dataset consists of 6,957 functions written in C, 2,244 in C++, 1,529 in C#, 3,187 in Go, 6,207 in Java, 5,066 in JavaScript, and 5,797 in Python. Our study encompasses all seven programming languages supported by REEF: C, C++, C#, Go, Java, JavaScript, and Python.
Data pre-processing. To adapt the REEF dataset for multilingual AVD tasks, our first step was to extract the vulnerable and non-vulnerable functions. We received commit data in raw format with patches. To analyze both the original and modified versions of vulnerable functions, we processed these patches using the Linux patch command. [^1] https://www.man7.org/linux/man-pages/man1/patch.1.html Following the commit data collection (Fan et al., 2020), we pre-processed the data by removing code comments to minimize bias. We then extracted function definition code using Tree-sitter (Brunsfeld, 2024), an efficient parser generator and incremental parsing library capable of analyzing multiple programming languages. To extract functions and their corresponding function-level and line-level labels, we iterated through function definitions in the pre-change file to match them with their post-change and pre-change function definitions. Specifically, the function names are used as identifiers to match and compare corresponding functions between the pre-change and post-change versions of the files. When multiple matching function definitions are found in the post-change file, we calculate the edit distance between the pre-change function and each post-change function. The pair with the smallest edit distance is then identified as the vulnerable and clean function pair. In other words, we define the pre-change function as the vulnerable function and the post-change function as the non-vulnerable function. Since many PLMs use the absolute position encoding, which is limited to the input length (e.g., 512), we filtered out functions that are greater than 512 in length. Finally, we obtained a total of 20,165 functions and corresponding labels. At the function-level, the dataset comprises 3,056 C, 1,792 C++, 427 C#, 2,905 Go, 3,235 Java, 5,468 JavaScript, and 3,282 Python functions. To obtain line-level labels, we first eliminated all non-vulnerable functions from the function-level dataset. We then applied DIFFLIB (Homepage, 2024) to match changed code lines by comparing each vulnerable function with its corresponding non-vulnerable version. DIFFLIB identifies three matching cases: (1) vulnerable functions fixed by adding code lines only, (2) vulnerable functions fixed by removing code lines only, and (3) vulnerable functions fixed by both removing and adding code lines. Since we need to locate specific line positions, only cases (2) and (3) can identify vulnerable line positions through deletion information. We use these two cases to obtain code line labels, i.e., line numbers and code lines, then filter out vulnerable functions where line labels cannot be extracted. Finally, we obtained a total of 8,134 vulnerable functions and the corresponding line labels. Finally, we obtained a total of 8,134 lines and corresponding labels. At the line-level, the dataset comprises 1,175 C, 651 C++, 151 C#, 919 Go, 1,375 Java, 2,496 JavaScript, and 1,367 Python functions.
Furthermore, to validate Tree-sitter’s robustness, we performed a sanity check. We randomly selected 377 samples, a sample ensuring a 95% confidence level with a 5% margin of error. The first author examined the function definitions between pre-change and post-change versions, verifying their consistency and alignment with the provided patch. Our manual inspection confirmed that Tree-sitter is reliable and generates no false positives.
Table 1. Statistical summary of the function-level multilingual vulnerability dataset
Languages #Training #Validation #Test #Total #Vul #Non-Vul C 2,444 305 307 3,056 1,541 1,515 C++ 1,432 179 181 1,792 911 881 C# 341 42 44 427 212 215 Go 2,323 290 292 2,905 1,462 1,443 Java 2,587 323 325 3,235 1,622 1,613 JavaScript 4,374 546 548 5,468 2,743 2,725 Python 2,625 328 329 3,282 1,642 1,640 Total 16,126 2,013 2,026 20,165 10,133 10,032
    #Training, #Validation, #Test, #Total, #Vul and #Non-Vul represent the number of training functions, validation functions, test functions, vulnerable functions, and non-vulnerable functions, respectively
Table 2. Statistical summary of the line-level multilingual vulnerability dataset
Languages #Training #Validation #Test #Total #Vul #Non-Vul C 937 118 120 1,175 3,764 32,191 C++ 528 63 60 651 2,223 14,380 C# 123 15 13 151 396 3,123 Go 735 92 92 919 3,123 19,958 Java 1,100 134 141 1,375 3,602 22,414 JavaScript 1,996 254 246 2,496 8,350 38,376 Python 1,094 135 138 1,367 3,702 21,165 Total 6,513 811 810 8,134 25,161 150,607
    #Training, #Validation, #Test, #Total, #Vul and #Non-Vul represent the number of training functions, validation functions, test functions, vulnerable lines, and non-vulnerable lines, respectively
Construction of training and test datasets. Similar to Fu and Tantithamthavorn (2022), we used random sampling to divide the REEF dataset of function-level and line-level multilingual vulnerability into training, validation, and test sets at an 8:1:1 ratio. To maintain consistent proportions of data across languages, we first allocated the data of each language according to the predetermined ratio. We then combined these portions to create the final dataset. As shown in Table 1, the function-level multilingual vulnerability dataset contains 16,126, 2,013, and 2,026 functions and corresponding labels for training, validation, and testing over seven languages. As shown in Table 2, the line-level multilingual vulnerability dataset contains 6,513, 811, and 810 vulnerable functions and corresponding line labels for training, validation, and testing across the studied languages.
Distribution of vulnerability severity. To analyze the dataset quality, we examined the severity levels of both function-level and line-level datasets, as shown in Table 1 and 2. We used the Common Vulnerability Scoring System (CVSS), a standardized method for measuring vulnerability severity. Using CVSS v4.0 Ratings, [^2] https://nvd.nist.gov/vuln-metrics/cvss we classified the vulnerabilities into four levels: low, medium, high, and critical. Figure 1 shows the severity distribution across the seven programming languages studied. In this multilingual vulnerability dataset, only 1.2% are rated as low severity, while 41.12% are classified as high severity and 25.09% as critical severity. These findings demonstrate that the real-world vulnerabilities provided by REEF are high-quality and predominantly target severe vulnerabilities across both function-level and line-level multilingual vulnerability.
Figure 1. The severity distribution of multilingual vulnerability dataset
3.2. Experimented Language Models
In our study, we investigated the performance of two types of language models in detecting multilingual vulnerability at both the function level and the line level: pre-trained language models and large language models. For the PLMs, we included six state-of-the-art PLMs for multilingual vulnerability detection, including:
  • CodeBERT (Feng et al., 2020) is a powerful pre-trained model that combines source code and natural language processing capabilities through its multilayer Transformer architecture.
  • CodeT5 (Wang et al., 2021) is a unified encoder-decoder model that integrates code token type information directly into its architecture. Building upon the T5 framework, it employs denoising sequence-to-sequence pre-training techniques.
  • CodeT5P (Wang et al., 2023a) extends CodeT5’s capabilities through an architecture featuring a shallow encoder paired with a deep decoder. The training process follows a multi-stage approach, beginning with unimodal training before progressing to bimodal data integration.
  • UniXCoder (Guo et al., 2022) leverages a unified cross-modal pre-trained architecture that incorporates multiple information sources like abstract syntax trees (AST) and code comments to enhance code representation, all built on transformer-based foundations.
  • LineVul (Fu and Tantithamthavorn, 2022) is a Transformer-based approach for predicting software vulnerabilities at the line level, aiming to enhance the granularity and accuracy of vulnerability detection. By leveraging the BERT architecture, LineVul captures long-term dependencies within code sequences, enabling precise identification of vulnerable lines.
  • Text-Embedding Models (OpenAI, 2024c) refer to a series of the latest generation of embedding models from OpenAI that offer enhanced representation capabilities for both text and code. Specifically, we utilized three text-embedding models: Text-Embedding-Ada-002, Text-Embedding-3-Small, and Text-Embedding-3-Large.
We focused exclusively on PLMs rather than employing GNNs or RNNs for multilingual vulnerability detection. This choice was motivated by PLMs’ proven strengths in vulnerability detection compared to RNNs and GNNs (Steenhoek et al., 2023). Moreover, alternative approaches (Zhang et al., 2023b; Liu et al., 2024; Steenhoek et al., 2024) heavily rely on Abstract Syntax Trees (ASTs). Since ASTs are inherently language-specific structures, creating a unified AST representation across different programming languages is impractical due to varying syntax and semantic constructs.
For LLMs, we studied five state-of-the-art models, including three open-source LLMs and two closed-source LLMs, which have demonstrated promising performance in various code-related tasks (Zhou et al., 2024a; Hou et al., 2023). The details of these LLMs are as follows:
  • DeepSeek-Coder (Guo et al., 2024) trained on a massive dataset of 2 trillion tokens spanning 87 programming languages. With its large 16K context window and specialized training in code completion tasks, it achieves leading performance among open-source models and even outperforms some closed-source solutions like Code Llama and GPT-3.5-Turbo in certain areas.
  • Code Llama (Roziere et al., 2023) derived from Llama 2, is a leading decoder-only language model specialized in code generation and infilling. The model achieves its capabilities through extensive fine-tuning on 500B tokens of code-focused data.
  • Llama 3 (Dubey et al., 2024) represents an innovative suite of large-scale multilingual language models leveraging the Transformer architecture to enhance performance across diverse natural language understanding tasks. Through sophisticated optimizations in data quality, training methodology, and architectural design, Llama 3 demonstrates marked advancements in natural language processing capabilities.
  • GPT-3.5-Turbo (OpenAI, 2022) is an improved version of GPT-3.5, optimized for conversations through reinforcement learning from human feedback. As one of the core models powering ChatGPT, it has become among the most widely used general-purpose language models.
  • GPT-4o (OpenAI, 2024b), introduced by OpenAI, is a multimodal AI model capable of processing and generating text, audio, and images in real time. It integrates these modalities within a unified architecture, enabling seamless interaction across different data types.
Similarly, we excluded LLM-based approaches like MSIVD (Yang et al., 2024) and GRACE (Lu et al., 2024) because they require AST construction, which is impractical for implementing unified ASTs across multiple programming languages and does not support line-level detection. We also omitted Vul-RAG (Du et al., 2024) because it cannot support or be extended to line-level vulnerability detection. As a result, we focused primarily on vanilla general and code-specific LLMs for multilingual vulnerability detection, following previous empirical studies (Fu et al., 2023a; Yin et al., 2024; Yin, 2024; Zhou et al., 2024c).
Moreover, we also introduce two types of dummy classifiers at the function-level and line-level detection as the ordinary baselines. The first type (${DummyClf}_{vul}$) predicts all functions or lines as vulnerable. The second type (${DummyClf}_{clean}$) predicts all functions or lines as clean. These baseline comparisons help us evaluate whether PLM and LLM’s performance on multilingual vulnerability detection differs meaningfully from extreme predictions, allowing us to assess whether PLM and LLM predictions show any systematic patterns rather than being purely random.
((a))
((b))
Figure 2. Zero-shot prompt template for multilingual vulnerability detection
3.3. Strategies for Large Language Models
We also investigated the effectiveness of LLMs’ various learning strategies for detecting multilingual vulnerability at both the function- and line-levels, since different learning strategies can significantly affect the LLMs’ performance. For instance, existing studies (Tian et al., 2024) have shown that fine-tuning strategies effectively enhance LLM performance by adapting general LLMs to specific tasks. Additionally, diverse prompting strategies have been introduced as plug-and-play solutions to enhance LLM performance in code-related tasks (Kojima et al., 2022). Thus, we devised three LLM strategies:
  • Zero-shot prompting strategy: We developed a prompt structure with system and user roles, following prior works (Zhou et al., 2024b; Zhang et al., 2023a). We directly employed a structured instruction and a vulnerable function to prompt LLMs for multilingual vulnerability detection at both function and line levels, without using examples. Figure 2 shows the zero-shot prompt templates for function-level and line-level multilingual vulnerability detection. For function-level detection, we specify that the input is a code function written in a particular programming language, then ask the LLM to determine whether the function is vulnerable or clean. For line-level detection, we specify that the input is a vulnerable function written in a particular programming language, then ask the LLM to identify the vulnerable line number and its corresponding code.
    ((a))
    ((b))
    Figure 3. Few-shot prompt template for multilingual vulnerability detection
  • Few-shot prompting strategy: It refers to providing LLMs with a small number of input-output examples in the prompt to help them understand the desired task before giving them a new input to complete. Specifically, it combines demonstration examples with a zero-shot prompt to form a new few-shot prompt, which is then fed to LLMs for detecting function-level or line-level multilingual vulnerability. Figure 3 presents the few-shot prompt template for function-level and line-level multilingual vulnerability detection. Based on prior research  (Pornprasit and Tantithamthavorn, 2024), when analyzing a given function, we select three demonstration examples from the REEF training dataset using BM25 (Robertson et al., 2009). These examples are used to prompt LLMs to detect function- and line-level multilingual vulnerabilities. BM25 was chosen as the sample selection method because previous studies (Gao et al., 2023; Yuan et al., 2023) demonstrate its superior performance over alternative approaches for software engineering tasks.
  • Instruction-tuning strategy: It enables LLMs to acquire task-specific knowledge and align with the desired response characteristics by supervised training on numerous instruction-filled function pairs. Ouyang et al. (2022) have shown that LLMs perform better on new, unfamiliar tasks when they are fine-tuned using diverse datasets that include natural language instructions.
    For function-level detection, an instruction-filled function pair contains three elements: a zero-shot prompt instruction (see Figure 2(a)), a given function, and a label indicating whether that function is vulnerable or clean. Using all 16,126 instances from the function-level training set, we constructed a function-level instruction-filled fine-tuning set. We then fine-tuned the LLMs on this fine-tuning set to detect function-level multilingual vulnerability using either zero-shot or few-shot prompts. Similarly, for line-level detection, an instruction-filled function pair consists of a zero-shot prompt instruction (see Figure 2(b)), a vulnerable function, and corresponding line labels. Using all 6,513 instances from the line-level training set, we created a line-level instruction-filled fine-tuning set. Then, we fine-tuned the LLMs on this set to detect line-level multilingual vulnerability using either zero-shot or few-shot prompts.
3.4. Metrics
Following existing work (Fu and Tantithamthavorn, 2022; Steenhoek et al., 2024), we adopted four widely-used metrics, Accuracy, Precision, Recall, and F1-score, to assess all our studied models’ effectiveness in detecting function-level multilingual vulnerability as a binary classification task. Accuracy measures the proportion of correctly predicted instances out of all instances. Precision quantifies how many of the predicted positive instances are actually correct, while Recall assesses the proportion of actual positive instances correctly predicted by the model. The F1-score combines precision and recall into a single metric, providing a balanced view especially useful when dealing with imbalanced datasets. Specifically, in our study, we used binary-averaged Precision, Recall, and F1-score because function-level vulnerability detection can be formulated as a binary classification task. Since the ratio between positive (vulnerable) and negative (non-vulnerable) instances in the function-level dataset is approximately 50:50 (see Table 1), we considered Accuracy a critical metric for measuring the models’ effectiveness in multilingual vulnerability detection.
For line-level multilingual vulnerability detection, we formulated the line-level detection task as a binary classification problem for each code line by following the existing work (Li et al., 2021b). Therefore, we also adopted Accuracy, Precision, Recall, and F1-score to assess all our studied models’ effectiveness in detecting line-level multilingual vulnerability.
Furthermore, we adopted the False Positive Rate (FPR) to measure how often a classifier incorrectly predicts the positive class when the actual class is negative, and False Negative Rate (FNR) to quantify how often the model mistakenly classifies positive instances as negative. In vulnerability detection, false positives (when clean code is flagged as vulnerable) cause developers to waste resources investigating secure code, while false negatives (when vulnerable code is flagged as clean) pose a fatal risk by leaving security flaws unaddressed. The trade-off between FPR and FNR, achieving lower FNR at the cost of lower FPR, is critical for measuring the intrinsic effectiveness of PLMs and LLMs in multilingual vulnerability detection. Specifically, FPR is defined as $\frac{FP}{FP+TN}$ where FP represents false positives and TN represents true negatives. FNR is defined as $\frac{FN}{FN+TP}$ where FN represents false negatives and TP represents true positives. A high FPR means many clean functions and lines remain unrecognized. A high FNR means many vulnerable functions or lines remain undetected.
3.5. Implementation and Environment
For pre-trained language models (PLMs), we replicated CodeBert, CodeT5, CodeT5P, UniXCoder, LineVul, and Text-Embedding-Models following existing works (Steenhoek et al., 2024). We obtained all open-source pre-trained checkpoints (i.e., CodeBert, CodeT5, CodeT5P, UniXCoder, and LineVul) from Huggingface (Wolf et al., 2019). Since Text-Embedding-Models are a series of closed-source PLMs, we accessed them through OpenAI’s API. For our experiments with closed-source LLMs, we accessed GPT-3.5-Turbo (model version gpt-3.5-turbo-0125) and GPT-4o (model version gpt-4o-2024-08-06) through OpenAI’s APIs (OpenAI, 2024a). For open-source LLMs, we utilized pre-trained checkpoints from Huggingface (Wolf et al., 2019) for DeepSeek-Coder (6.7B parameters), Code Llama (7B parameters), and Llama 3 (8B parameters). We implemented Low-Rank Adaptation (LoRA (Hu et al., 2021)) during fine-tuning to optimize computational efficiency and prevent overfitting. We conducted all the experiments on an Intel Xeon CPU Gold-6342 machine with 512 GB RAM, Ubuntu 20.04.6, and two A800 GPUs. More implementation details can be found in our replication package. To facilitate future research, we have made all the used datasets and model execution scripts publicly available.
4. Evaluation Results
4.1. RQ1: How effective are PLMs and LLMs in detecting multilingual vulnerabilities at the function-level?
Approach. This research question provides a comparative analysis of performance between various PLMs and LLMs with different strategies in detecting function-level multilingual vulnerability. Specifically, we investigated the effectiveness of eight PLMs and five LLMs. The model details are summarized in Section 3.2. We now detail the training/inference process below.
Regarding PLMs, they are pre-trained on a corpus of natural language and source code snippets from various programming languages, but they have not been exposed to the task of function-level vulnerability detection, particularly across multilingual vulnerability. Prior studies (Fu and Tantithamthavorn, 2022; Steenhoek et al., 2023, 2024) have demonstrated the effectiveness of fine-tuned PLMs for function-level vulnerability detection. Therefore, we fine-tune all studied open-source PLMs (i.e., CodeBert, CodeT5, CodeT5P, UniXCoder, and LineVul) on the function-level training set. The model architecture for function-level multilingual vulnerability detection combines a PLM’s encoder with a downstream dense layer that acts as a binary classifier. We train all parameters of both the encoder and binary classifier through supervised learning on the function-level training set. The trained encoder outputs a code embedding representing the given input given code. This embedding is then fed into the trained binary classifier, which produces a binomial distribution indicating the probability of the given function being vulnerable. For closed-source PLMs (i.e., Text-Embedding-Ada-002, Text-Embedding-3-Small, and Text-Embedding-3-Large), we follow the same training and inference approach. Since these PLMs do not allow access to their parameters, we treat them as frozen encoders and only train the binary classifier.
Regarding LLMs, they are decoder-only models with billions of parameters, pre-trained on vast data corpora of text and source code to process and generate human-like language. Although LLMs have not been trained specifically for vulnerability detection, we can leverage their powerful in-context learning capabilities and pre-trained prior knowledge of various programming languages to perform function-level multilingual vulnerability detection with or without instruction tuning. Without instruction tuning, we directly applied zero-shot or few-shot prompts to have LLMs predict either 0 or 1, where 1 indicates a vulnerable function and 0 indicates a non-vulnerable one. With instruction tuning, we first trained the LLMs through supervised fine-tuning on the instruction-filled function-level training set. Then, we applied zero-shot or few-shot prompts to obtain the predictions of trained LLMs.
To evaluate the effectiveness of the studied PLMs and LLMs in detecting function-level multilingual vulnerabilities, we employed multiple metrics, including Accuracy, Precision, Recall, F1-score, FPR, and FNR as detailed in Section 3.4. Given the balanced distribution between positive and negative labels, Accuracy was used as the primary evaluation metric.
Table 3. Performance of PLMs and LLMs for function-level multilingual vulnerability detection
Techniques Accuracy Recall Precision F1-score FPR FNR Dummy Classifiers ${DummyClf}_{vul}$ 0.5030 1.0000 0.5030 0.6693 1.0000 0.0000 ${DummyClf}_{clean}$ 0.4970 0.0000 0.0000 0.0000 0.0000 1.0000 PLMs Text-Embedding-3-Large 0.5064 0.9696 0.5049 0.6640 0.9623 0.0304 Text-Embedding-3-Small 0.5044 0.9676 0.5038 0.6626 0.9643 0.0324 Text-Embedding-Ada-002 0.5030 1.0000 0.5030 0.6693 1.0000 0.0000 CodeBERT 0.5173 1.0000 0.5103 0.6757 0.9712 0.0000 LineVul 0.5173 1.0000 0.5103 0.6757 0.9712 0.0000 UniXcoder 0.5814 0.8930 0.5518 0.6822 0.7339 0.1070 CodeT5 0.5854 0.9342 0.5519 0.6939 0.7676 0.0658 CodeT5P 0.6037 0.9529 0.5626 0.7075 0.7496 0.0471 LLMs (Zero-Shot Prompting) DeepSeek-Coder 0.5005 0.1796 0.5097 0.2656 0.1748 0.8204 Code Llama 0.4877 0.9156 0.4950 0.6426 0.9454 0.0844 Llama 3 0.5005 0.5015 0.5034 0.5025 0.5005 0.4985 GPT-3.5-Turbo 0.4827 0.6183 0.4888 0.5459 0.6544 0.3817 GPT-4o 0.4985 0.3631 0.5020 0.4214 0.2632 0.7399 LLMs (Few-Shot Prompting) DeepSeek-Coder 0.3870 0.5309 0.4146 0.4656 0.7587 0.4691 Code Llama 0.4763 0.6251 0.4840 0.5456 0.6743 0.3749 Llama 3 0.4299 0.5260 0.4437 0.4814 0.6673 0.4740 GPT-3.5-Turbo 0.3618 0.4426 0.3835 0.4109 0.7200 0.5574 GPT-4o 0.5035 0.4347 0.5074 0.4683 0.4270 0.5653 LLMs (Instruction Tuning + Zero-Shot Prompting) DeepSeek-Coder 0.5039 0.3023 0.5116 0.3800 0.2920 0.6977 Code Llama 0.4783 0.5103 0.4824 0.4959 0.5531 0.4897 Llama 3 0.5192 0.5348 0.5215 0.5281 0.4965 0.4652 GPT-3.5-Turbo 0.5454 0.5132 0.5517 0.5318 0.4220 0.4868 GPT-4o 0.5874 0.4622 0.6206 0.5298 0.2859 0.5378 LLMs (Instruction Tuning + Few-Shot Prompting) DeepSeek-Coder 0.4906 0.4789 0.4934 0.4861 0.4975 0.5201 Code Llama 0.5094 0.6408 0.5098 0.5678 0.6226 0.3592 Llama 3 0.4946 0.4848 0.4975 0.4911 0.4955 0.5152 GPT-3.5-Turbo 0.5000 0.4524 0.5033 0.4765 0.4518 0.5475 GPT-4o 0.7196 0.6722 0.7454 0.7069 0.2324 0.3278
Results. Table 3 shows the comparison results among PLMs and LLMs in terms of average Accuracy, Recall, Precision, F1-score, FPR, and FNR in detecting function-level multilingual vulnerability. Within each metric, the values in bold indicate the model that exhibits the best performance among all PLMs and LLMs. Figure 4 presents the performance of PLMs and LLMs with instruction tuning and few-shot prompting in seven programming languages in terms of Accuracy.
Observation 1: Among PLMs, CodeT5P exhibits superior performance in function-level multilingual vulnerability detection. In the PLMs’ category, text-embedding models like Text-Embedding-3-Large, Text-Embedding-3-Small, and Text-Embedding-Ada-002 all hover around 0.5000 in Accuracy, indicating near-random performance on a balanced dataset. Despite high Recall scores (all >0.9600), their low Precision (around 0.5000) results in modest F1-scores (around 0.6600). This suggests a bias toward over-identifying functions as vulnerable, in other words, a high FPR (all >0.9600). Similarly, CodeBERT and LineVul achieve slightly higher Accuracy (0.5163 and 0.5173) and perfect Recall (1.0000) but suffer from low Precision (0.5098 and 0.5103). This leads to middling F1-scores (0.6753 and 0.6757) and limited practical utility, as they cannot reliably discriminate between vulnerable and non-vulnerable functions.
More modern code-specific models show improved balance. UniXcoder and CodeT5 both outperform general text-embedding models and simple code-specific models, with UniXcoder reaching 0.5814 Accuracy and CodeT5 achieving 0.5854. These models strike a better trade-off between Recall and Precision, resulting in higher F1-scores (0.6822 and 0.6939, respectively). Most notably, CodeT5P delivers the strongest performance across all PLMs, achieving the highest Accuracy of 0.6037 along with a solid F1-score of 0.7075. Furthermore, the CodeT5P reaches FPR of 0.7496 and FNR of 0.0471, which implies that although CodeT5P can detect more vulnerable functions, it comes at the cost of wrongly detecting more clean functions. This indicates that CodeT5P, which incorporates identifier-aware training objectives and enriched representation learning, is particularly effective at capturing the multilingual semantic nuances required for function-level multilingual vulnerability detection. Overall, the results demonstrate that while traditional PLMs struggle with precision and overall Accuracy, task-aware models like CodeT5P offer significant improvements, making them more suitable for multilingual vulnerability detection scenarios at the function-level.
Observation 2: LLMs exhibit limitations in detecting function-level multilingual vulnerability through zero-shot or few-shot prompting strategies alone. In evaluating the effectiveness of LLMs for function-level multilingual vulnerability detection under zero-shot and few-shot prompting settings, Accuracy emerges as a critical metric due to the balanced nature of the dataset. Under zero-shot prompting setting, the average Accuracies of LLMs across seven programming languages were 0.5005 for DeepSeek-Coder, 0.4877 for Code Llama, 0.5005 for Llama 3, 0.4758 for GPT-3.5-Turbo, and 0.4985 for GPT-4. Specifically, DeepSeek-Coder achieves the highest Accuracy (0.5005), closely followed by Llama 3 (0.5005) and Code Llama (0.4877). Although DeepSeek-Coder exhibits the highest Accuracy, it achieves the lowest Recall (0.1796), leading to a poor F1-score (0.2656). This suggests a tendency to overpredict non-vulnerabilities. In contrast, Code Llama stands out with the highest Recall (0.9156). However, its relatively lower Precision (0.4950) results in a moderate F1-score (0.6426), indicating a tendency to overpredict vulnerabilities, leading to a higher FPR (0.9454) GPT-3.5-Turbo and GPT-4o demonstrate lower Recall but more balanced Precision-Recall trade-offs, though with modest Accuracy scores (0.4827 and 0.4985, respectively).
Under the few-shot prompting setting, this strategy does not consistently improve accuracy. In fact, most LLMs show decreased performance compared to zero-shot prompting, with GPT-4o being the only exception. Specifically, DeepSeek-Coder drops sharply to 0.3870 Accuracy, Code Llama to 0.4763, Llama 3 to 0.4299, and GPT-3.5-Turbo to 0.3618, indicating that providing few examples may confuse rather than help the models. In this case, GPT-4o is the only exception, slightly improving to 0.5035 Accuracy, though its Recall (0.4347) and F1-score (0.4683) remain relatively low. These results highlight that zero-shot prompting yields more stable Accuracy for LLMs, and few-shot prompting does not guarantee improvement. Overall, LLMs using only zero-shot or few-shot prompting show limited effectiveness in function-level multilingual vulnerability detection, emphasizing the need for further adaptation to the task domain.
((a))
((b))
Figure 4. Performance of PLMs and LLMs across seven programming languages in function-level multilingual vulnerability detection (y-axis: Accuracy)
Observation 3: Instruction tuning is critical to improve the LLM’s effectiveness in detecting function-level multilingual vulnerability. Instruction tuning is able to enhance the function-level multilingual vulnerability detection capabilities of LLMs, especially when paired with prompting strategies. In the instruction tuning with zero-shot prompting setting, GPT-4o achieves the highest Accuracy (0.5874), outperforming GPT-3.5-Turbo (0.5454), Code Llama (0.4783), Llama 3 (0.5192), and DeepSeek-Coder (0.5039). Notably, most of the instruction-tuned LLMs show improvement in Accuracy, Precision, and F1-scores, compared to their non-instruction-tuned counterparts. For instance, the most significant accuracy improvements were seen in GPT-4o (17.83%) and GPT-3.5-Turbo (14.63%). These results suggest that instruction tuning enables better generalization and contextual understanding of multilingual vulnerability patterns, even without in-context examples.
When few-shot prompting was combined with instruction tuning, most LLMs exhibited the same tendency as with few-shot prompting alone, i.e., LLMs’ performance deteriorated, except for GPT-4o. In this case, GPT-4o achieves a remarkable Accuracy of 0.7196, the highest across all LLMs with different strategies, along with the highest Precision (0.7454) and F1-score (0.7069). Additionally, GPT-4o achieves a relatively balanced ratio between its FNR of 0.3278 and FPR of 0.2324, indicating its ability to effectively detect vulnerable functions while maintaining accurate predictions for clean functions. This suggests that GPT-4 has strong discriminative capabilities and robustness in detecting function-level multilingual vulnerability. Compared to the best-performing PLM, CodeT5P, GPT-4o exhibits a significant advantage in Accuracy and Precision, while maintaining a competitive F1-score. Although CodeT5P shows higher Recall (0.9529 vs. 0.6722), CodeT5P’s FPR (0.7496) is also significantly higher than GPT-4o’s FPR (0.2324). In contrast, GPT-4o’s balanced trade-off across all metrics, particularly its much lower harmonic mean between FPR and FNR ($\frac{0.2324+0.3278}{2}$ < $\frac{0.7496+0.0471}{2}$), making it more practical in real-world applications. This comparison underscores the growing potential of instruction-tuned LLMs, especially GPT-4o, as powerful and adaptable techniques for function-level multilingual vulnerability detection when paired with minimal task-specific examples.
Observation 4: The effectiveness of GPT-4o with instruction tuning and few-shot prompting is promising on seven different programming languages, and superior to other PLMs and LLMs. As shown in Figure 4(a), the three embedding models, Text-Embedding-3-Large, Text-Embedding-3-Small, and Text-Embedding-Ada-002, demonstrate consistent performance levels, maintaining an approximate Accuracy of 0.5000 across all seven programming languages evaluated. The evaluation reveals that CodeBERT and LineVul demonstrate limited effectiveness, with an approximate Accuracy score of 0.5000 when detecting function-level vulnerability on C, C#, C++, JavaScript, Java, and Python. Notably, their performance improves in the context of the Go language, where both models achieve an Accuracy of 0.5959. In contrast, UniXcoder, CodeT5, and CodeT5P consistently outperform other PLMs in Accuracy measurements across seven programming languages. Specifically, CodeT5P stands out as the best-performing PLM, demonstrating robust performance with an Accuracy of 0.5081 in C, 0.6136 in C#, 0.5249 in C++, 0.6062 in Go, 0.7062 in JavaScript, 0.6031 in Java, and 0.5623 in Python. UniXcoder and CodeT5 stand out as suboptimal PLMs, showing relatively weak performance with Accuracies of 0.4919 and 0.5081 in C, 0.5455 and 0.5455 in C#, 0.5193 and 0.5028 in C++, 0.6473 and 0.6130 in Go, 0.6204 and 0.6898 in JavaScript, 0.6031 and 0.5169 in Java, and 0.5593 and 0.5775 in Python.
Figure 4(b) shifts the focus to LLMs, revealing GPT-4o as the standout model in terms of Accuracy across all seven programming languages. Specifically, GPT-4o achieves the strongest performance with an Accuracy of 0.7557 in C, 0.7273 in C#, 0.7901 in C++, 0.8082 in Go, 0.6642 in JavaScript, 0.7169 in Java, and 0.6626 in Python. Furthermore, GPT-4o not only surpasses CodeT5P in high-level languages but also exhibits higher and more stable Accuracy across traditionally challenging languages such as Go, C, and C++. In several cases (e.g., Python, JavaScript), GPT-4o exceeds 0.6500 Accuracy, reflecting its superior capacity to understand diverse syntactic structures and vulnerability patterns across languages. This comparison underscores the strength of instruction-tuned LLMs with few-shot prompting, positioning GPT-4o as a more versatile and reliable model for function-level multilingual vulnerability detection.
RQ1 Summary: GPT-4o (with instruction tuning and few-shot prompting) and CodeT5P achieved the highest performance among LLMs and PLMs, scoring 0.7196 and 0.6037 in average Accuracy across seven programming languages, respectively. Furthermore, GPT-4o with instruction tuning and few-shot prompting outperformed all other PLMs and LLMs across all seven studied languages, showing improvements of 19.20% and 85.94% in average Accuracy. In evaluating performance across multiple programming languages using instruction tuning and few-shot prompting, GPT-4o reached the highest accuracy of 0.8082 with Go and the lowest accuracy of 0.6626 with Python.
4.2. RQ2: How effective are PLMs and LLMs in detecting multilingual vulnerabilities at the line level?
Approach. This research question focuses on a comparative analysis of performance between different PLMs and LLMs with various strategies in detecting line-level multilingual vulnerability. Specifically, we investigated the effectiveness of four PLMs and five LLMs. In this case, we excluded three Text-Embedding Models and LineVul. The Text-Embedding Models can only provide a single embedding for a given function, making them unsuitable for line-level detection. LineVul uses an unsupervised learning approach, which would be unfair to compare with the supervised learning method used in our experiment. We detail the training and inference process below.
Regarding PLMs, we adopt a line-level embedding approach to align with the task of detecting vulnerabilities at the line-level. Specifically, we represented each vulnerable function using up to $n_{\text{lines}}=113$ lines, with each line containing up to $n_{\text{tokens}}=31$ tokens. This design is based on the observation that 95% of functions in our dataset contain fewer than 113 lines. Each input function is embedded into a tensor of shape $(n_{\text{lines}},n_{\text{tokens}},d_{\text{model}})$, where $d_{\text{model}}=768$ is the hidden dimension of the embedding. To obtain line-level representations, we summarized each line by applying a single-layer Gated Recurrent Unit (GRU) over the token dimension ($n_{\text{tokens}}$), producing a tensor of shape $(n_{\text{lines}},d_{\text{model}})$. We formulated the task as a multi-label classification problem, aiming to identify which lines within a vulnerable function are vulnerable. Each vulnerable function contains at least one, and potentially multiple, vulnerable lines. We added a classification head for each PLM that outputs a single value per line, followed by a sigmoid activation function. A line is classified as vulnerable if its output exceeds a threshold of 0.5.
Regarding LLMs, we expected to use their in-context learning capabilities to achieve line-level multilingual vulnerability detection with or without instruction tuning. Line-level detection required LLMs to identify vulnerable code lines, but previous studies (Fu et al., 2023b) and our initial experiments have shown that LLMs cannot accurately distinguish code lines based on line breaks. To address this issue, we re-structured the functions by adding tags (i.e., Line N:) in front of each code line. Without instruction tuning, we used zero-shot or few-shot prompts to guide LLMs in predicting vulnerable line numbers (i.e., tags of lines) and their corresponding code. With instruction tuning, we first performed supervised fine-tuning on our instruction-filled line-level training set. We then prompted these instruction-tuned LLMs to generate vulnerable line numbers and corresponding code using zero-shot or few-shot prompting.
To evaluate the effectiveness of the studied PLMs and LLMs in detecting line-level multilingual vulnerability, we used common classification metrics: Accuracy, Precision, Recall, F1-score, FPR, and FNR, as detailed in 3.4. Since our task involves classifying each line of code within a function, we faced a challenge with imbalanced data, i.e., most code lines are non-vulnerable, creating an uneven distribution in our line-level dataset. For this reason, we chose F1-score as our primary metric.
Table 4. Performance of PLMs and LLMs for line-level multilingual vulnerability detection
Techniques Accuracy Recall Precision F1-score FPR FNR Dummy Classifier ${DummyClf}_{vul}$ 0.0267 1.0000 0.0267 0.0520 1.0000 0.0000 ${DummyClf}_{clean}$ 0.9733 0.0000 0.0000 0.0000 0.0000 1.0000 PLMs CodeBERT 0.9203 0.8822 0.2353 0.3715 0.0787 0.1177 UniXcoder 0.8880 0.8879 0.1787 0.2975 0.1120 0.1121 CodeT5 0.9472 0.8119 0.3123 0.4511 0.0491 0.1881 CodeT5P 0.9563 0.7673 0.3536 0.4841 0.0385 0.2327 LLMs (Zero-Shot Prompting) DeepSeek-Coder 0.9536 0.3002 0.2246 0.2570 0.0284 0.6997 Code Llama 0.9589 0.2237 0.2272 0.2254 0.0208 0.7763 Llama 3 0.9307 0.4847 0.1891 0.2720 0.0570 0.5153 GPT-3.5-Turbo 0.9557 0.3215 0.2475 0.2797 0.0229 0.8393 GPT-4o 0.9537 0.4286 0.2700 0.3313 0.0318 0.5714 LLMs (Few-Shot Prompting) DeepSeek-Coder 0.9521 0.3493 0.2344 0.2805 0.0313 0.6507 Code Llama 0.9666 0.3018 0.3536 0.3257 0.0151 0.6982 Llama 3 0.8980 0.7509 0.1738 0.2822 0.0980 0.2491 GPT-3.5-Turbo 0.9774 0.4372 0.6057 0.5078 0.0078 0.5628 GPT-4o 0.9802 0.5689 0.6461 0.6050 0.0086 0.4311 LLMs (Instruction Tuning + Zero-Shot Prompting) DeepSeek-Coder 0.9404 0.2155 0.1297 0.1619 0.0397 0.7845 Code Llama 0.8771 0.5902 0.1234 0.2042 0.1150 0.4098 Llama 3 0.9415 0.3403 0.1819 0.2371 0.0420 0.6597 GPT-3.5-Turbo 0.9578 0.5186 0.3211 0.3966 0.0301 0.4814 GPT-4o 0.9729 0.4241 0.4910 0.4551 0.0121 0.5759 LLMs (Instruction Tuning + Few-Shot Prompting) DeepSeek-Coder 0.7149 0.2851 0.0278 0.0507 0.2733 0.7149 Code Llama 0.8688 0.6221 0.1207 0.2022 0.1243 0.3779 Llama 3 0.9220 0.4720 0.1647 0.2442 0.0657 0.5280 GPT-3.5-Turbo 0.9716 0.5804 0.4735 0.5215 0.0177 0.4196 GPT-4o 0.9830 0.6307 0.7012 0.6641 0.0074 0.3693
Results. Table 4 shows the comparison results among PLMs and LLMs in terms of average Accuracy, Recall, Precision, F1-score, FPR, and FNR in detecting line-level multilingual vulnerability. Within each metric, the values in bold indicate the model that exhibits the best performance among all PLMs and LLMs. Figure 5 presents the performance of PLMs, LLMs with zero-shot prompting, and LLMs with instruction tuning and few-shot prompting in seven programming languages in terms of F1-score.
Observation 5: CodeT5P is the best-performing model in identifying line-level multilingual vulnerability among PLMs. Among the evaluated PLMs, CodeT5P demonstrates the strongest overall performance in detecting line-level vulnerabilities, achieving the highest F1-score of 0.4841. This comparatively high F1-score primarily stems from its balanced approach, reflected in a moderate Recall of 0.7673 and Precision of 0.3536. Additionally, CodeT5P balanced the trade-off between predicting false positives and false negatives, achieving the lowest FPR (0.0385) among studied PLMs. Although UniXCoder achieves higher Recall (0.8879), indicating sensitivity in identifying vulnerable lines, its significantly lower Precision (0.1787) and higher FPR (0.1120) severely impacts its overall performance, yielding a much lower F1-score of 0.2975. CodeBert exhibits the modest performance among PLMs, with an F1-score of 0.3715, despite having relatively high Recall (0.8822). Thus, among traditional PLMs, CodeT5P emerges as the most viable choice for line-level vulnerability detection, effectively balancing FPR and FNR to mitigate the challenges of the imbalanced dataset.
Observation 6: While using prompting alone, few-shot prompting can improve the effectiveness of LLMs in detecting line-level multilingual vulnerability. Under zero-shot prompting, all evaluated LLMs show limited practical effectiveness despite achieving notably high Accuracy scores. Code Llama attains the highest Accuracy (0.9589); however, it exhibits an extremely low F1-score (0.2254), reflecting significant weaknesses in Precision (0.2272), Recall (0.2237), and FNR (0.7763). Similarly, models like GPT-4o and GPT-3.5-Turbo display relatively high Accuracies (0.9537 and 0.9557, respectively) but similarly low F1-scores (0.3313 and 0.2797) and high FNR (0.5714 and 0.8393). The consistently low Recall/Precision and high FNR across zero-shot prompting indicate a clear inability of LLMs to effectively detect vulnerable lines without further guidance or contextual clues, underscoring the insufficiency of using LLMs in isolation under zero-shot conditions for multilingual line-level vulnerability detection.
Few-shot prompting improves LLMs’ performance through in-context examples. GPT-4o demonstrates the best performance, surpassing all other PLMs with a significantly higher F1-score of 0.6050, supported by strong Precision (0.6461) and moderate Recall (0.5689). Compared to the zero-shot prompting setting, GPT-4o’s FPR significantly decreases from 0.0318 to 0.0086, while its FNR also reduces from 0.5714 to 0.4311. GPT-3.5-Turbo also shows notable improvement, with its F1-score rising to 0.5078. This improvement is primarily driven by enhanced Precision (0.6057), although Recall remains limited (0.4372). These results demonstrate that few-shot prompting alone, without instruction tuning, can partially address the issue of imbalanced labels (where vulnerable lines are fewer than non-vulnerable lines) to improve LLMs’ performance in detecting multilingual vulnerabilities at the line-level.
((a))
((b))
Figure 5. Performance of PLMs and LLMs across seven programming languages in line-level multilingual vulnerability detection (y-axis: F1-score)
Observation 7: Instruction tuning is also beneficial for improving LLMs in detecting multilingual vulnerabilities at the line-level. Finally, the addition of instruction tuning combined with few-shot prompting significantly enhances LLM performance, clearly outperforming all other prompting strategies. GPT-4o, in particular, achieves the highest overall performance with an impressive F1-score of 0.6641. This high F1-score results from a robust Precision of 0.7012 and substantial Recall of 0.6307, effectively handling the challenge posed by dataset imbalance. Additionally, GPT-4o reaches the lowest FPR of 0.0074 and the lowest FNR of 0.3693 among the studied LLMs (except for Llama 3 with few-shot prompting). This indicates an excellent balance between correctly identifying vulnerable and clean lines. GPT-3.5-Turbo also shows notable improvements under this condition, achieving an F1-score of 0.5215, demonstrating the broader positive impact of instruction tuning combined with few-shot prompting. While using instruction tuning with few-shot prompting, GPT-4o and GPT-3.5-Turbo significantly outperform the dummy classifier in F1-score and Precision. This demonstrates that their predictions are substantially more accurate than random guesses for line-level detection.
When comparing GPT-4o with the best-performing PLM, CodeT5P (F1-score: 0.4841), GPT-4o’s superiority becomes even more evident. GPT-4o notably surpasses CodeT5P not only in F1-score (0.6641 vs. 0.4841) but also significantly in Precision (0.7012 vs. 0.3536). Although CodeT5P achieves slightly higher Recall (0.7673 vs. GPT-4o’s 0.6307), its lower Precision substantially limits its practical utility, as it generates many more false-positive predictions. Supporting evidence comes from the FPR and FNR metrics. GPT-4o achieves a much lower FPR (0.0074 vs. CodeT5P’s 0.0385), though it has a higher FNR (0.3693 vs. CodeT5P’s 0.2327).
In terms of F1-score on each programming language, GPT-4o from Figure 5 demonstrates a clear superiority in detecting line-level vulnerabilities. GPT-4o consistently achieves higher F1-scores across all languages, particularly excelling in Go and JavaScript, whereas CodeT5P shows more moderate performance with notable gaps in languages like C, Go, and Java. Specifically, GPT-4o demonstrates the best performance with F1-scores of 0.5478 in C, 0.4348 in C#, 0.5774 in C++, 0.7078 in Go, 0.7815 in JavaScript, 0.6391 in Java, and 0.6055 in Python. These results are consistently superior to CodeT5P (except in C#), which achieved F1-scores of 0.3727 in C, 0.4750 in C#, 0.4768 in C++, 0.4146 in Go, 0.6556 in JavaScript, 0.4167 in Java, and 0.4519 in Python. Consequently, these results indicate that GPT-4o’s instruction tuning combined with few-shot prompting substantially enhances its ability to generalize across diverse programming languages for line-level vulnerability detection, significantly surpassing the capability of traditional PLMs like CodeT5P.
RQ2 Summary: In a comparison of LLMs and PLMs, GPT-4o with instruction tuning and few-shot prompting emerged as the top performer, achieving an F1-score of 0.6641 across seven programming languages, followed by CodeT5P at 0.4841. Additionally, GPT-4o demonstrated superior performance across all languages, with F1-score improvements ranging from 9.77% to 310.19% compared to other PLMs and LLMs. In evaluating performance across multiple programming languages using instruction tuning and few-shot prompting, GPT-4o achieves the highest F1-score of 0.7815 with JavaScript and the lowest F1-score of 0.4348 with C#.
4.3. RQ3: What are the strengths and weaknesses of the PLMs and LLMs in multilingual vulnerability detection?
Approach. This research question aims to deepen our understanding of the performance of various PLMs and LLMs with different strategies in detecting multilingual vulnerability at the function-level and line-level. According to RQ1 and RQ2, we selected the best-performing PLMs and LLMs with various strategies as our representative models. In our function-level multilingual vulnerability detection analysis, we selected models based on Accuracy metric. We chose CodeT5P for PLMs and evaluated four LLM variants: Llama 3 using zero-shot prompting (ZSP), and three GPT-4o configurations - few-shot prompting (FSP), instruction tuning with zero-shot prompting (ITZSP), and instruction tuning with few-shot prompting (ITFSP). For line-level multilingual vulnerability detection, we based our model selection on F1-score performance. We maintained CodeT5P as our PLM representative and utilized the four GPT-4o variants (ZSP, FSP, ITZSP, and ITFSP) for LLM representatives. Our orthogonality analysis consists of the following three perspectives:
  • The unique correct and incorrect detection. Certain PLMs and LLMs may be prone to detecting more vulnerabilities. Therefore, we further explored the orthogonality between the studied PLMs and LLMs by using Venn diagrams to assess the unique correct detection and incorrect detection. To generate the Venn diagrams, we first categorized correct/incorrect detections by PLMs and LLMs into distinct sets. We then compared these sets to analyze their overlap and unique characteristics. This analysis revealed which correct/incorrect detections were made by PLMs alone, by LLMs alone, and by both models together.
  • The tendency in predicting the Top-25 dangerous CWE-IDs. Since certain PLMs and LLMs tend to be more effective at detecting specific CWE types, we further evaluated the accuracy of detecting specific CWE types at both function-level and line-level. We used test data from the 2023 CWE Top 25 Most Dangerous Software Weaknesses, published by the CWE community. [^3] https://cwe.mitre.org/top25/archive/2023/2023_top25_list.html Note that our test data encompasses all of the top 25 CWE-IDs.
  • The effectiveness across the different vulnerability severity. For the last perspective, PLMs and LLMs tend to detect different vulnerability severity, which can affect the effectiveness of models in real-world scenarios. Hence, we evaluated the number of correct detections at both function- and line-levels using test data categorized by CVSS v4.0 ratings.
Results. Figures 6 and 7 show Venn diagrams depicting the intersection of correct and incorrect detections at both function-level and line-level among the studied PLMs and LLMs. The overlapping regions indicate shared correct or incorrect detections, while the non-overlapping areas show detections unique correct or incorrect predictions to each model. Figures 8 and 9 illustrate the performance comparison of multilingual vulnerability detection at both function- and line-level, with results categorized by vulnerability severity. Tables 5 and 6 further present the detection performance of studied PLMs and LLMs for the top 25 most dangerous CWE-IDs in 2023 at function-level and line-level. The values highlighted in bold denote the model or strategy that exhibits the best performance for the corresponding CWE-ID.
Figure 6. The unique correct/incorrect detection of function-level multilingual vulnerability in PLMs and LLMs
Figure 7. The unique correct/incorrect detection of line-level multilingual vulnerability in PLMs and LLMs
Observation 8: In terms of unique correct and incorrect detection, GPT-4o with ITFSP outperforms the other studied models, achieving better results in both function-level detection (with fewer incorrect detections) and line-level detection (with more unique correct detections and zero incorrect ones). For function-level detection, Figure  6a demonstrates that CodeT5P achieves superior performance in unique correct detections compared to LLMs using the four studied strategies. CodeT5P identifies 108 unique correct detections, substantially outperforming Llama 3 with ZSP (48), GPT-4o with FSP (14), GPT-4o with ITZSP (24), and GPT-4o with ITFSP (28). However, Figure  6b shows that GPT-4o with ITFSP has fewer unique incorrect detections at the function-level, only 11 compared to CodeT5P (136), Llama 3 with ZSP (189), GPT-4o with FSP (85), and GPT-4o with ITZSP (68). For line-level detection, Figure  7a shows that GPT-4o with ITFSP outperforms all studied PLMs and LLMs, with 56 unique correct detections compared to CodeT5P (6), GPT-4o with ZSP (2), GPT-4o with FSP (20), and GPT-4o with ITZSP (15). Figure 7b demonstrates that GPT-4o with ITFSP achieves the best performance in unique incorrect detections. GPT-4o with ITFSP has only 0 unique incorrect detections, substantially outperforming CodeT5P (13), GPT-4o with ZSP (101), GPT-4o with FSP (2), and GPT-4o with ITZSP (1). These results further demonstrate the effectiveness of GPT-4o with ITFSP in both function- and line-level multilingual vulnerability detection, reinforcing our findings from RQ1 and RQ2.
Table 5. The percentage of correct detection for function-level multilingual vulnerability detection using two representative techniques across the Top 25 most dangerous CWE-IDs in 2023
Ranking CWE-ID Llama 3 (ZSP) GPT-4o (FSP) GPT-4o (ITZSP) GPT-4o (ITFSP) CodeT5P Total 1 CWE-787 (Out-of-bounds Write) 19(46.34%) 21(51.22%) 22(53.66%) 34(82.93%) 23(56.1%) 41 2 CWE-79 (Cross-site Scripting) 97(48.99%) 115(58.08%) 126(63.64%) 155(78.28%) 125(63.13%) 198 3 CWE-89 (SQL Injection) 21(45.65%) 33(71.74%) 33(71.74%) 36(78.26%) 27(58.7%) 46 4 CWE-416 (Use After Free) 20(58.82%) 17(50.0%) 17(50.0%) 26(76.47%) 19(55.88%) 34 5 CWE-78 (OS Command Injection) 5(33.33%) 8(53.33%) 10(66.67%) 12(80.0%) 5(33.33%) 15 6 CWE-20 (Improper Input Validation) 43(46.74%) 46(50.0%) 47(51.09%) 62(67.39%) 52(56.52%) 92 7 CWE-125 (Out-of-bounds Read) 28(54.9%) 21(41.18%) 24(47.06%) 35(68.63%) 31(60.78%) 51 8 CWE-22 (Path Traversal) 45(59.21%) 43(56.58%) 56(73.68%) 60(78.95%) 43(56.58%) 76 9 CWE-352 (Cross-Site Request Forgery) 42(52.5%) 56(70.0%) 56(70.0%) 62(77.5%) 62(77.5%) 80 10 CWE-434 (Unrestricted Upload of File with Dangerous Type) 2(40.0%) 3(60.0%) 3(60.0%) 3(60.0%) 3(60.0%) 5 11 CWE-862 (Missing Authorization) 1(100.0%) 0(0.0%) 0(0.0%) 1(100.0%) 1(100.0%) 1 12 CWE-476 (NULL Pointer Dereference) 19(42.22%) 28(62.22%) 23(51.11%) 34(75.56%) 19(42.22%) 45 13 CWE-287 (Improper Authentication) 64(45.71%) 82(58.57%) 105(75.0%) 108(77.14%) 110(78.57%) 140 14 CWE-190 (Integer Overflow or Wraparound) 17(53.12%) 18(56.25%) 18(56.25%) 27(84.38%) 18(56.25%) 32 15 CWE-502 (Deserialization of Untrusted Data) 18(62.07%) 17(58.62%) 16(55.17%) 21(72.41%) 18(62.07%) 29 16 CWE-77 (Command Injection) 13(50.0%) 15(57.69%) 19(73.08%) 18(69.23%) 14(53.85%) 26 17 CWE-119 (Improper Restriction of Operations within the Bounds of a Memory Buffer) 27(43.55%) 28(45.16%) 33(53.23%) 49(79.03%) 30(48.39%) 62 18 CWE-798 (Use of Hard-coded Credentials) 2(66.67%) 1(33.33%) 1(33.33%) 2(66.67%) 1(33.33%) 3 19 CWE-918 (Server-Side Request Forgery) 13(39.39%) 13(39.39%) 16(48.48%) 22(66.67%) 20(60.61%) 33 20 CWE-306 (Missing Authentication for Critical Function) 9(50.0%) 15(83.33%) 13(72.22%) 15(83.33%) 9(50.0%) 18 21 CWE-362 (Race Condition) 7(53.85%) 6(46.15%) 7(53.85%) 10(76.92%) 5(38.46%) 13 22 CWE-269 (Improper Privilege Management) 7(63.64%) 1(9.09%) 3(27.27%) 8(72.73%) 8(72.73%) 11 23 CWE-94 (Code Injection) 11(45.83%) 13(54.17%) 15(62.5%) 17(70.83%) 18(75.0%) 24 24 CWE-863 (Incorrect Authorization) 28(65.12%) 18(41.86%) 24(55.81%) 32(74.42%) 22(51.16%) 43 25 CWE-276 (Incorrect Default Permissions) 1(100.0%) 1(100.0%) 0(0.0%) 0(0.0%) 1(100.0%) 1 Mean 559(49.96%) 619(55.32%) 687(61.39%) 849(75.87%) 684(61.13%) 1,119 * The numbers in parentheses represent the percentage of correct prediction of the corresponding techniques, while the "Total" column represents the total amount of data corresponding to CWE-ID in the test dataset of this paper
Table 6. The percentage of correct detection for line-level multilingual vulnerability detection using two representative techniques across the Top 25 most dangerous CWE-IDs in 2023
Ranking CWE-ID GPT-4o (ZSP) GPT-4o (FSP) GPT-4o (ITZSP) GPT-4o (ITFSP) CodeT5P Total 1 CWE-787 (Out-of-bounds Write) 0(0.0%) 0(0.0%) 1(6.25%) 1(6.25%) 0(0.0%) 16 2 CWE-79 (Cross-site Scripting) 9(10.59%) 46(54.12%) 44(51.76%) 57(67.06%) 26(30.59%) 85 3 CWE-89 (SQL Injection) 5(27.78%) 5(27.78%) 9(50.0%) 9(50.0%) 5(27.78%) 18 4 CWE-416 (Use After Free) 1(6.25%) 5(31.25%) 5(31.25%) 7(43.75%) 2(12.5%) 16 5 CWE-78 (OS Command Injection) 1(20.0%) 3(60.0%) 1(20.0%) 4(80.0%) 2(40.0%) 5 6 CWE-20 (Improper Input Validation) 2(4.88%) 16(39.02%) 10(24.39%) 17(41.46%) 3(7.32%) 41 7 CWE-125 (Out-of-bounds Read) 0(0.0%) 1(4.55%) 3(13.64%) 6(27.27%) 1(4.55%) 22 8 CWE-22 (Path Traversal) 0(0.0%) 9(32.14%) 4(14.29%) 9(32.14%) 2(7.14%) 28 9 CWE-352 (Cross-Site Request Forgery) 13(33.33%) 23(58.97%) 20(51.28%) 24(61.54%) 17(43.59%) 39 10 CWE-434 (Unrestricted Upload of File with Dangerous Type) 0(0.0%) 0(0.0%) 0(0.0%) 0(0.0%) 0(0.0%) 1 11 CWE-862 (Missing Authorization) 0(0.0%) 0(0.0%) 0(0.0%) 0(0.0%) 0(0.0%) 1 12 CWE-476 (NULL Pointer Dereference) 2(14.29%) 6(42.86%) 3(21.43%) 9(64.29%) 1(7.14%) 14 13 CWE-287 (Improper Authentication) 0(0.0%) 19(51.35%) 19(51.35%) 22(59.46%) 14(37.84%) 37 14 CWE-190 (Integer Overflow or Wraparound) 1(10.0%) 3(30.0%) 3(30.0%) 5(50.0%) 2(20.0%) 10 15 CWE-502 (Deserialization of Untrusted Data) 2(15.38%) 6(46.15%) 7(53.85%) 8(61.54%) 3(23.08%) 13 16 CWE-77 (Command Injection) 2(16.67%) 7(58.33%) 7(58.33%) 10(83.33%) 4(33.33%) 12 17 CWE-119 (Improper Restriction of Operations within the Bounds of a Memory Buffer) 0(0.0%) 5(21.74%) 1(4.35%) 5(21.74%) 1(4.35%) 23 18 CWE-798 (Use of Hard-coded Credentials) 0(0.0%) 1(100.0%) 0(0.0%) 1(100.0%) 0(0.0%) 1 19 CWE-918 (Server-Side Request Forgery) 2(14.29%) 6(42.86%) 4(28.57%) 6(42.86%) 3(21.43%) 14 20 CWE-306 (Missing Authentication for Critical Function) 0(0.0%) 5(55.56%) 2(22.22%) 6(66.67%) 4(44.44%) 9 21 CWE-362 (Race Condition) 0(0.0%) 0(0.0%) 0(0.0%) 2(50.0%) 1(25.0%) 4 22 CWE-269 (Improper Privilege Management) 0(0.0%) 1(25.0%) 0(0.0%) 1(25.0%) 0(0.0%) 4 23 CWE-94 (Code Injection) 2(16.67%) 7(58.33%) 6(50.0%) 7(58.33%) 3(25.0%) 12 24 CWE-863 (Incorrect Authorization) 2(13.33%) 5(33.33%) 4(26.67%) 7(46.67%) 2(13.33%) 15 25 CWE-276 (Incorrect Default Permissions) 0(0.0%) 0(0.0%) 0(0.0%) 1(100.0%) 0(0.0%) 1 0 Mean 44(9.98%) 179(40.59%) 153(34.69%) 224(50.79%) 96(21.77%) 441 * The numbers in parentheses represent the percentage of correct prediction of the corresponding techniques, while the "Total" column represents the total amount of data corresponding to the CWE-ID in the test dataset of this paper
Observation 9: Among the Top 25 most dangerous vulnerabilities, such as CWE-89 (SQL Injection), GPT-4o with ITFSP demonstrates superior detection capabilities in both function-level and line-level detection, achieving remarkable detection rates of 75.87% and 50.79%, respectively. Regarding function-level detection, Table  5 shows that GPT-4o with ITFSP outperforms all studied PLMs and LLMs. Specifically, GPT-4o with ITFSP successfully detects 75.87% of vulnerable and clean functions associated with the top 25 most dangerous CWE-IDs in 2023, substantially outperforming Llama 3 with ZSP (49.96%), GPT-4o with FSP (55.32%), GPT-4o with ITZSP (61.39%), and CodeT5P (61.13%). GPT-4o with ITFSP achieves the best performance on 23 out of 25 CWE-IDs, while the second-best model, CodeT5P, leads in only 7. Regarding line-level detection, Table  6 shows the same trend as function-level detection, where GPT-4o with ITFSP is superior to other studied PLMs and LLMs. GPT-4o with ITFSP successfully detects 50.79% of vulnerable and clean functions for the top 25 most dangerous CWE-IDs in 2023, significantly outperforming GPT-4o with ZSP (9.98%), GPT-4o with FSP (40.59%), GPT-4o with ITZSP (34.69%), and CodeT5P (21.77%). These findings highlight the effectiveness of GPT-4o with ITFSP in addressing dangerous vulnerabilities at both function-level and line-level, reinforcing the results from RQ1 and RQ2.
Figure 8. The number of correct detections based on CVSS severity at the function-level in PLMs and LLMs
Figure 9. The number of correct detections based on CVSS severity at the line-level in PLMs and LLMs
Observation 10: Across all CVSS severity levels, GPT-4o with ITFSP demonstrates superior capability in identifying multilingual vulnerabilities, showing exceptional performance at both function and line-level detection compared to other studied models. For function-level detection, Figure 8 demonstrates that GPT-4o with ITFSP performs better at detecting High, Medium, and Low severity vulnerabilities. However, GPT-4o with ITFSP performs slightly worse than CodeT5P when detecting Critical severity issues. For example, GPT-4o with ITFSP exhibits 630 detections of High severity vulnerability, surpassing 514 by CodeT5P, 438 by Llama 3 with ZSP, 465 by GPT-4o with FSP, and 518 by GPT-4o with ITZSP. GPT-4o with ITFSP exhibits 305 detections of Critical severity vulnerability, which is slightly lower than 321 by CodeT5P. For line-level detection, Figure 9 shows that GPT-4o with ITFSP outperforms all studied PLMs and LLMs at detecting Critical, High, Medium, and Low severity vulnerabilities. For example, GPT-4o with ITFSP exhibits 134 detections of Critical severity vulnerability, significantly surpassing 76 by CodeT5P, 14 by GPT-4o with (ZSP), 117 by GPT-4o with FSP, and 111 by GPT-4o with ITZSP. These findings underscore that GPT-4o with ITFSP consistently delivers superior performance in correctly detecting vulnerabilities at both function-level and line-level granularity across different CVSS severity categories, reflecting the advantage of incorporating instruction tuning and few-shot prompting in enhancing LLMs.
RQ3 Summary: For both function-level and line-level analysis, GPT-4o, with instruction tuning and few-shot prompting, outperforms the PLM representative (CodeT5P) and other studied LLM representatives. This superiority is shown in unique correct/incorrect multilingual detection, vulnerability detection in the Top-25 dangerous CWE-IDs, and vulnerability detection across various severity levels.
5. Discussion
5.1. How Does the Model Size Affect the LLM’s Strategy?
Table 7. Performance of Larger LLMs in detecting function-level multilingual vulnerability
Techniques Accuracy Recall Precision F1-Score FPR FNR LLMs (Zero-Shot Prompting) Code Llama (7B) 0.4877 0.9156 0.4950 0.6426 0.9454 0.0844 Code Llama (70B) 0.5089 0.7360 0.5081 0.6012 0.7210 0.2640 Llama 3 (8B) 0.5005 0.5015 0.5034 0.5025 0.5005 0.4985 Llama 3 (70B) 0.4956 0.7733 0.4991 0.6066 0.7855 0.2267 LLMs (Few-Shot Prompting) Code Llama (7B) 0.4763 0.6251 0.4840 0.5456 0.6743 0.3749 Code Llama (70B) 0.4931 0.6133 0.4968 0.5490 0.6286 0.3867 Llama 3 (8B) 0.4299 0.5260 0.4437 0.4814 0.6673 0.4740 Llama 3 (70B) 0.4659 0.6762 0.4781 0.5602 0.7468 0.3238 LLMs (Instruction Tuning + Zero-Shot Prompting) Code Llama (7B) 0.4783 0.5103 0.4824 0.4959 0.5531 0.4897 Code Llama (70B) 0.5035 0.5339 0.5060 0.5196 0.5273 0.4661 Llama 3 (8B) 0.5192 0.5348 0.5215 0.5281 0.4965 0.4652 Llama 3 (70B) 0.5079 0.6300 0.5087 0.5629 0.6157 0.3700 LLMs (Instruction Tuning + Few-Shot Prompting) Code Llama (7B) 0.5094 0.6408 0.5098 0.5678 0.6226 0.3592 Code Llama (70B) 0.4941 0.5564 0.4973 0.5252 0.5690 0.4436 Llama 3 (8B) 0.4946 0.4848 0.4975 0.4911 0.4955 0.5152 Llama 3 (70B) 0.5093 0.5996 0.5104 0.5514 0.5819 0.4004
Table 8. Performance of Larger LLMs in detecting line-level multilingual vulnerability
Techniques Accuracy Recall Precision F1-Score FPR FNR LLMs (Zero-Shot Prompting) Code Llama (7B) 0.9589 0.2237 0.2272 0.2254 0.0208 0.7763 Code Llama (70B) 0.9628 0.1321 0.2016 0.1596 0.0144 0.8679 Llama 3 (8B) 0.9307 0.4847 0.1891 0.2720 0.0570 0.5153 Llama 3 (70B) 0.9619 0.3092 0.2960 0.3025 0.0202 0.6908 LLMs (Few-Shot Prompting) Code Llama (7B) 0.9666 0.3018 0.3536 0.3257 0.0151 0.6982 Code Llama (70B) 0.9565 0.2385 0.2161 0.2267 0.0237 0.7616 Llama 3 (8B) 0.8980 0.7509 0.1738 0.2822 0.0980 0.2491 Llama 3 (70B) 0.9745 0.5190 0.5237 0.5214 0.0130 0.4810 LLMs (Instruction Tuning + Zero-Shot Prompting) Code Llama (7B) 0.8771 0.5902 0.1234 0.2042 0.1150 0.4098 Code Llama (70B) 0.9565 0.2568 0.2251 0.2400 0.0243 0.7431 Llama 3 (8B) 0.9415 0.3403 0.1819 0.2371 0.0420 0.6597 Llama 3 (70B) 0.9611 0.2573 0.2654 0.2613 0.0195 0.7427 LLMs (Instruction Tuning + Few-Shot Prompting) Code Llama (7B) 0.8688 0.6221 0.1207 0.2022 0.1243 0.3779 Code Llama (70B) 0.9496 0.3918 0.2345 0.2934 0.0351 0.6082 Llama 3 (8B) 0.9220 0.4720 0.1647 0.2442 0.0657 0.5280 Llama 3 (70B) 0.9752 0.5092 0.5378 0.5231 0.0120 0.4908
Our prior experiments demonstrate that combining instruction tuning with zero-shot/few-shot prompting does not effectively improve all LLMs. This limitation likely stems from the LLMs’ basic capabilities, which typically improve with larger model sizes. Based on performance results from RQ1 and RQ2, we selected Code Llama (70B) and Llama 3 (70B) to investigate how model size influences different LLM strategies and to minimize potential bias in our findings. The implementation details for both Code Llama (70B) and Llama 3 (70B) remain consistent with Section 3. As in RQ1 and RQ2, due to differences in positive and negative sample distribution, we evaluate using Accuracy for the function-level task and F1-Score for the line-level task.
As shown in Table 7, Code Llama (70B) and Llama 3 (70B) achieve Accuracies of 0.5089 and 0.4956, respectively, in detecting function-level vulnerability using zero-shot prompting. The Code Llama (70B) shows slight improvement compared to Code Llama (7B), while Llama 3 (70B) performs slightly worse than Llama 3 (8B). With instruction tuning and zero-shot prompting, both Code Llama (70B) and Llama 3 (70B) demonstrate similar performance patterns. When using few-shot prompting, Code Llama (70B) and Llama 3 (70B) achieve Accuracies of 0.4931 and 0.4659, exceeding the performance of their smaller counterparts, Code Llama (7B) at 0.4763 and Llama 3 (8B) at 0.4299. When combining instruction tuning with few-shot prompting, Code Llama (70B) and Llama 3 (70B) perform similarly to their smaller versions, achieving Accuracies of 0.4941 and 0.5093 compared to 0.5094 and 0.4946, respectively. As shown in Table 8, for line-level detection, Code Llama (70B) reaches F1-Scores of 0.1596 and 0.2267 when using zero-shot prompting and few-shot prompting, which is degraded from 0.2254 and 0.3257 by Code Llama (7B). Llama 3 (70B) is superior to its smaller version while using zero-shot and few-shot prompting, with F1-Score of 0.3025 and 0.5214 compared to 0.2720 and 0.2822, respectively. Instruction tuning can better improve the larger LLMs at detecting line-level vulnerability. For example, when using instruction tuning with few-shot prompting, Code Llama (70B) and Llama 3 (70B) reach F1-Scores of 0.2934 and 0.5231, significantly superior to 0.2022 by Code Llama (7B) and 0.2442 by Llama 3 (8B).
These results indicate that larger LLMs are not the decisive factor in function-level detection performance. While few-shot prompting works better with larger models, instruction tuning can lead to under-fitting problems when there are too many parameters. Consequently, when using a larger model for function-level detection, careful instruction tuning and appropriate hyper-parameter (e.g. number of adapters in LoRA) selection are essential. Larger LLMs have a more significant impact on line-level vulnerability detection, especially when using Llama 3, likely due to their more concentrated training sets and more informative labels. The differences between these models, despite having the same parameters, can be attributed to the models’ inherent characteristics. For example, Llama 3 differs from Code Llama by using a new tokenizer and training on a larger token dataset, which can significantly affect the LLMs’ fundamental ability and then further affect the effectiveness of instruction tuning.
5.2. Comparison between Reasoning LLMs and Non-Reasoning LLMs
Table 9. Comparison between reasoning and non-reasoning LLMs.
Techniques Accuracy Recall Precision F1-Score FPR FNR Non-reasoning LLMs (Function-level) DeepSeek-Coder 0.5005 0.1796 0.5097 0.2656 0.1748 0.8204 Code Llama 0.4877 0.9156 0.4950 0.6426 0.9454 0.0844 Llama 3 0.5005 0.5015 0.5034 0.5025 0.5005 0.4985 GPT-3.5-Turbo 0.4827 0.6183 0.4888 0.5459 0.6544 0.3817 GPT-4o 0.4985 0.3631 0.5020 0.4214 0.2632 0.7399 Reasoning LLMs (Function-level) DeepSeek-R1 0.5217 0.6585 0.5193 0.5807 0.6167 0.3415 QwQ-plus 0.5346 0.4838 0.5418 0.5111 0.4141 0.5162 Non-reasoning LLMs (Line-level) DeepSeek-Coder 0.9536 0.3002 0.2246 0.2570 0.0284 0.6997 Code Llama 0.9589 0.2237 0.2272 0.2254 0.0208 0.7763 Llama 3 0.9307 0.4847 0.1891 0.2720 0.0570 0.5153 GPT-3.5-Turbo 0.9557 0.3215 0.2475 0.2797 0.0229 0.8393 GPT-4o 0.9537 0.4286 0.2700 0.3313 0.0318 0.5714 Reasoning LLMs (Line-level) DeepSeek-R1 0.9692 0.2184 0.3698 0.2746 0.0102 0.7816 QwQ-plus 0.9655 0.2953 0.3344 0.3136 0.0161 0.7047
With the development of LLMs, a new category called reasoning LLMs (Treude and Kula, 2025) has emerged to handle more complex tasks. The distinction between reasoning and non-reasoning LLMs primarily lies in their approach to problem-solving and the complexity of tasks they can handle. Reasoning LLMs can perform step-by-step processing and tackle multi-step tasks like code-related challenges. For example, DeepSeek-R1 (Guo et al., 2025) and QwQ-plus (Cloud, 2024) use reinforcement learning techniques (e.g., GRPO) to enhance their reasoning abilities. Non-reasoning LLMs, such as GPT-4o, generate responses based on learned patterns without explicit intermediate reasoning steps. To further validate our findings, we selected DeepSeek-R1 and QwQ-plus as our studied reasoning LLMs to evaluate their effectiveness in detecting function-level and line-level multilingual vulnerabilities. To ensure experimental fairness and maximize the capabilities of reasoning LLMs, we selected the DeepSeek-R1 with 671B parameters and the default QwQ-plus model, whose parameters are not publicly available. To minimize prompt-related variability, we used the same zero-shot prompt described in Section 3 for both reasoning and non-reasoning LLMs.
Table 9 compares how reasoning and non-reasoning LLMs perform at detecting multilingual vulnerabilities at both function-level and line-level. For function-level detection, reasoning LLMs outperform all studied non-reasoning LLMs. DeepSeek-R1 and QwQ-plus achieve Accuracies of 0.5217 and 0.5346, respectively, significantly higher than DeepSeek-Coder (0.5005), Code Llama (0.4877), Llama 3 (0.5005), GPT-3.5-Turbo (0.4758), and GPT-4o (0.4985). For line-level detection, reasoning LLMs perform similarly to non-reasoning LLMs in terms of F1-score. For instance, DeepSeek-R1 achieves 0.2746 and QwQ-plus reaches 0.3136, comparable to GPT-3.5-Turbo’s 0.2797 and GPT-4o’s 0.3313.
Although reasoning LLMs are superior to non-reasoning LLMs at detecting function-level vulnerabilities, their effectiveness is only slightly better than random guessing, and their performance is not significantly better than non-reasoning LLMs. When using reasoning LLMs to detect line-level vulnerabilities, the improvements are minimal and match the performance of non-reasoning LLMs. Since reasoning LLMs require longer inference times but offer minimal or no improvements, implementing them for multilingual vulnerability detection requires careful balancing of efficiency and effectiveness. While we have explored the feasibility of reasoning LLMs for multilingual vulnerability detection, the effectiveness of fine-tuning these models using reinforcement learning still needs to be explored.
6. Threats to Validity
External Threats primarily stem from our choice of multilingual vulnerability dataset and automated vulnerability detection approaches. Our study focused on the REEF dataset covering seven programming languages, which may limit the generalizability of our findings to other languages. To address this limitation, we plan to expand our vulnerability collection to include more programming languages and timelines using the REEF collection framework. To minimize methodological threats, we performed a comprehensive literature review to ensure our selected PLMs and LLMs for automated vulnerability detection represent current state-of-the-art solutions.
Internal Threat relates to our implementation and utilization of the studied approaches. To address this, we carefully implemented all existing approaches following their original replication packages, with two authors conducting thorough code reviews. For PLMs, we followed the hyperparameter settings recommended by the original authors to ensure a fair reproduction of their approaches. However, the 512-token input limit of the smaller PLMs we used, which cannot be extended during fine-tuning, may restrict the models’ ability to capture context in longer code snippets. We acknowledge this architectural constraint as a limitation of our experimental setup. For LLMs, we utilized official channels, accessing public models (DeepSeek-Coder, Code Llama, and Llama3) via HuggingFace and invoking APIs (GPT-3.5-Turbo and GPT-4o) according to documented guidelines.
Construct Threats arise from our construction of vulnerable functions, clean functions, and metrics for evaluating PLMs and LLMs. We used Tree-sitter to parse commit data and collect before-change and after-change functions, which introduces a potential risk of incorrect function identification. To mitigate this threat, we conducted sanity checks on randomly selected samples to verify the tool’s robustness. For evaluation, we employed established metrics from function-level and line-level vulnerability detection, including Accuracy, Precision, Recall, and F1-score.
7. Conclusion
This work systematically evaluated the performance of pre-trained language models (PLMs) and large language models (LLMs) across seven programming languages for multilingual vulnerability detection at both function and line granularity. Our findings highlight that LLMs, particularly GPT-4o enhanced with instruction tuning and few-shot prompting, significantly outperform existing PLMs like CodeT5P. Specifically, GPT-4o achieved superior accuracy in function-level detection and markedly higher precision and F1-scores in line-level detection tasks, demonstrating its effectiveness and versatility in multilingual contexts. Additionally, orthogonality analysis revealed that GPT-4o not only detects vulnerabilities uniquely missed by other models but also excels in identifying high-severity vulnerabilities, emphasizing its practical value in real-world security applications. Despite these promising results, our study also reveals that the size and reasoning capabilities of LLMs do not necessarily correlate directly with improved vulnerability detection performance.
Meanwhile, our study opens up several potential future research directions, including: extending the scale of multilingual vulnerability benchmark, introducing different intermediate representations for unifying the code representation, and refining the PLM/LLM learning strategies. For instance, beyond supervised fine-tuning and instruction tuning, strategies like multi-task learning or meta-learning could better align PLMs and LLMs with multilingual domains by reducing the performance gap between high-resource and low-resource programming languages. In general, our work provides critical insight and foundational knowledge that can guide the advancement of robust multilingual vulnerability detection tools, addressing key limitations identified in existing approaches.
Acknowledgements.
This work was supported by National Key Research and Development Program of China (Grant No. 2024YFB4506300), the National Natural Science Foundation of China (Grant No. 62322208), JST under the Adopting Sustainable Partnerships for Innovative Research Ecosystem (ASPIRE) program, Grant Number JPMJAP2415, JSPS for the KAKENHI grants (JP21H04877, JP22K18630), Bilateral Program grant JPJSBP120239929, and the Inamori Research Institute for Science for supporting Yasutaka Kamei via the InaRIS Fellowship.